# auxmulticols[1]

### v1

@@ 4. Companies’ perspective
(C6.1) Usability challenges. The process for users to exercise data rights is generally considered to have poor usability. Consider privacy policies, for instance. Numerous studies have concluded that these policies are typically too verbose and complex for the average person, as evidenced by various readability metrics, including word count and Flesch scores (128; 108; 219; 10; 156; 143; 106; 236; 182; 191; 85; 18; 223; 114). On average, these policies contain about 2000 words and are written at a 13th-grade reading level, presenting substantial comprehension challenges for many users. Consistent with this, users frequently express frustration with the difficulty of understanding privacy policies (128; 175; 16; 129). The implementation requirements according to the GDPR data rights can degrade the usability of platforms (232; 223).
@@
These usability issues, along with dark patterns that result in discouraging users’ exercise of data rights, extend beyond the right to information and impact other rights such as the rights to access, erasure, data portability, and objection (177; 121; 158; 82; 139; 85; 84; 226; 169; 224; 83; 248; 55). A platform allows users to access their data only when they email them in a specific format (224). Numerous studies also point to the difficulty of understanding downloaded data due to its volume, disorganization, use of jargon, and a lack of description (226; 172; 169; 248). Additionally, research reveals that users often struggle to find ways to delete their personal data from apps and services (177; 121; 158). For example, disproportionate amounts of time to find a “delete" (158) or “opt-out” (84) button or the need to call the service team to do so (121). Indeed, the absence of straightforward account deletion options is a pervasive issue in online services, affecting about 77% of them (82). Overall, the evidence suggests that exercising data rights is a nontrivial task for many users.
@@
(C6.2) Usability Successes. Despite the above challenges, there are instances where users have found the process of exercising their rights straightforward, particularly in the cases of the rights to access and erasure (7; 135; 226; 84). For example, a survey shows that 58% of users who had previously requested companies to erase their personal information found the process straightforward (135). Furthermore, a study found that users had no difficulty in requesting a copy of their personal information from a loyalty card provider (7). Certain websites, like Amazon, Facebook, and Google, provide online portal tools to facilitate the rights request process, eliminating the need for users to contact the companies directly (226). Consequently, the usability of exercising data rights varies significantly across different applications and data rights types.
@@
(C6.3) High-usability Data Rights Tools. Given the usability issues associated with exercising data rights, the development of user-friendly data rights tools has been a natural response. Developers and researchers proposed and launched many tools and services that allow users to easily make use of data rights (187; 129; 15; 27; 196; 178; 20; 148; 112; 111; 38; 169; 168). While several studies validate that such tools can increase the effectiveness of data rights (187; 129; 69; 15; 27; 196; 178; 38), we can also find some testimonies that raise doubts about their effects (219; 69; 189; 112; 7; 34; 169). Given this, the effects of “high usability” tools are still unclear.
@@
(C6.3.1) Effectiveness. Empirical research suggests that high-usability data rights tools can effectively assist users in asserting their data rights. Emphasis is on tools that aim to mitigate the common challenge of understanding privacy policies (187; 129; 15; 27; 196; 178; 20; 148; 112; 38; 168). They commonly employ visualization, summarization, and user engagement strategies to simplify data processing comprehension.
@@
For instance, designs using visualization techniques display icons representing collected data or diagrams illustrating data flows, aiding intuitive understanding (196; 178; 27). In contrast, some tools prioritize brevity and highlight crucial points in privacy policies, presenting users with a summarized version (129; 15; 241; 168). Users concur that these techniques–visualization and summarization–facilitate understanding of policy content compared to conventional privacy policies. Moreover, research suggests the significant role of color in enhancing comprehension, with vibrant hues being particularly effective. Such approaches could also be applied to improve the readability of data obtained via the right to access (226; 169).
@@
Another category of tools promotes user engagement as a strategy to increase awareness of policy content (112; 111). This is predicated on the established theory that active learning fosters more effective comprehension (33; 167; 231). Empirical evidence confirms enhanced user awareness when users engage actively with policy content via “Drag-and-Drop” and “Question-and-Answer” features. User experiments also reveal a higher likelihood or willingness of data rights assertion when interacting with a chatbot (38; 147). In conclusion, ample evidence supports the efficacy of designs employing visualization, summarization, or user engagement in improving the effectiveness of data rights.
@@
Interestingly, many users do not expect companies to provide easy-to-use online interfaces to exercise data rights. In an interview, while some users, particularly those more interested in their data rights, express a desire for online data rights tools, the majority view traditional means such as email or phone as the most appropriate channels to claim their access rights (7). Some users even express a preference for offline data rights exercises, favoring direct interaction with a responsible party (34; 153). Furthermore, an evaluation of a purportedly “high-usability” data dashboard (228) revealed unfavorable responses. A subset of participants demonstrated no inclination or necessity to utilize it, some voiced objections to the very notion of “high-usability” dashboards, and others dismissed data management as inconsequential.
@@
The implementation of data rights should not expose users to additional threats. Current practices, however, often fail to satisfy this criterion. Specifically, research indicates a risk of leaking personal information to unauthorized third parties when users exercise their right to access (39; 40; 59; 32). Companies must verify the identities of those making data rights requests to ensure they are the actual data owners. A flawed identification process can result in the leakage of sensitive personal information to unauthorized individuals. Several studies reveal system vulnerabilities susceptible to impersonation via counterfeit emails (39; 40; 59). These studies demonstrate how easily an attacker can deceive service providers by using an email address similar to the original one, with success rates ranging from approximately 30% to 70% (39; 40; 59). Some systems do not even implement email encryption, despite requiring users to submit identification documents via email (39; 32). Hence, severe threats often arise when users exercise their right to access.
@@
(C8.1) Efficient data use. Several studies contend that the implementation of GDPR data rights significantly enhances companies’ data processing awareness, leading to improved user data management and increased competency in utilizing data (92; 135; 61; 171; 25; 199). The growing responsibility of companies in the information economy can also foster better decision-making and risk assessment (61). Large companies, in particular, seem to strongly recognize the benefits of improved data management processes (171).
@@
However, to assess whether data rights lead to efficient data use, it is important to consider not only companies’ data processing but also users’ data sharing behaviors. If data rights cause users to be more reticent about sharing their data, it could compromise the efficiency of data use by companies, despite improvements in internal data management processes. As discussed in U4.2, the evidence on this matter is conflicting (35; 146; 42; 160; 21). While some studies suggest that empowering users may generally incentivize them to disclose more personal data (35; 146; 42; 21), one report indicates that explicitly linking user data rights with the term GDPR can have the opposite effect (160). Therefore, in contrast to previous studies, we posit that it remains uncertain whether GDPR data rights enhance data use efficiency within companies. This area warrants further research.
@@
(C8.2) Positive corporate image. There is evidence of an increased trust of users in companies that foster data rights (135; 237; 144; 90). Survey results suggest that service providers can bolster their credibility by offering options to modify/delete personal information and maintaining transparency about data processing (135; 144; 90). A separate survey indicates that U.S. consumers place greater trust in companies that voluntarily extend GDPR data rights to non-European users (237). These findings suggest that user empowerment can enhance a company’s image.
@@
However, the relationship between data rights and trust is multifaceted. The extent to which data rights can rehabilitate a company’s image, if it is initially negative, remains uncertain. Under conditions of low trust, users may question the data rights themselves (176; 16; 52; 42; 165), making it challenging for companies to anticipate the positive effects of endorsing data rights. For instance, some interviewees express skepticism about whether companies would actually delete their personal information upon request (52). Another study reveals a perception that companies deliberately obfuscate their privacy policies to discourage user engagement (16). Some users also mistrust the information presented by companies (219).
@@
Low trust in companies appears to be a widespread issue among users. Research shows that, even when users express satisfaction with privacy laws like GDPR aimed at empowering them, they tend to lack confidence in companies’ compliance with such laws (7; 135; 52; 165). Consequently, trust may not increase substantially even when companies implement data rights, given the overall low level of trust in companies.
@@
Moreover, even if data rights significantly enhance trust in companies, this could have a counter-effect: as users trust service providers more, their willingness to exercise data rights may decrease, as discussed in U2.2 (108; 187). In summary, the relationship between data rights and trust is complex and warrants further exploration.
@@ 5. Regulators’ perspective
European regulators adopt a myriad of strategies to gauge and understand the effectiveness of GDPR data rights. Primarily, they interact with citizens through the complaints they receive, serving as a valuable resource for assessing user awareness of data rights (100; 150; 109; 131; 49). The France Data Protection Authority (DPA) reported a 64% increase in complaints in 2019 compared to 2018, interpreting this as an indication that EU citizens have “strongly embraced the GDPR” (109; 131). However, as discussed in U1, this assertion may not hold true. If the volume of complaints was low prior to the enforcement of the GDPR, a 64% increase may not necessarily imply a substantial rise in user awareness and engagement with GDPR data rights.
@@
Regulators also recognize the importance of evaluating the GDPR (179; 64). Evidence suggests that regulators understand the challenges associated with implementing data rights. Among the most cited difficulties is designing GDPR-compliant ML systems (50; 98; 170). Certain regulators have also broached the challenges of implementing GDPR data rights within blockchain systems (48). They actively research ML and blockchain technology and foster discussions through conferences and seminars (50; 75; 205; 48).
Moreover, legal experts, alongside regulators, highlight the difficulties in implementing data rights, specifically the right to erasure (132; 173). In interviews, UK data protection and privacy law experts expressed mixed views on the practicality of the right to erasure (132).
They also express concerns over certain websites deciding not to collect personal information from their users post-GDPR, arguing that it degrades the online user experience and stems from a misunderstanding of the GDPR (173). However, in spite of the challenges and unintended consequences, regulators and legal professionals largely agree that the GDPR has strengthened individual empowerment (132; 64).
@@
Post-GDPR, regulators have enacted supplementary measures to enhance users’ data rights. For instance, their websites feature numerous articles and posters promoting the concept of GDPR data rights to users and companies (45; 96; 29; 46; 3).
@@
Beyond these, regulators have proposed new guidelines and devised strategies to address various issues relating to online marketing, dark patterns, and cookies (76; 97; 49). As a result, regulators have been proactive in formulating new guidelines and strategizing to enhance user data rights. However, these guidelines also underscore the substantial costs companies may incur in complying with the GDPR.
@@
Ensuring robust enforcement is paramount to the protection of data rights, encompassing the detection and penalization of corporations infringing upon these rights via fines and other sanctions. Despite substantial endeavors by regulators to mitigate the challenges encountered by users and companies (refer to R1 and R2), ample evidence underscores the current inadequacy and deficiency of GDPR data rights enforcement. (102; 200; 185; 17; 238; 54; 6; 51).
A multitude of studies examining court cases have exposed trends in GDPR enforcement (185; 17; 6; 238; 54). These analyses reveal that imposed fines have typically been marginal in relation to the comprehensive economic and societal repercussions of GDPR noncompliance. Interestingly, even DPAs have adopted a conservative stance, refraining from levying fines to their maximum potential.
@@
The uneven enforcement across Europe can significantly prolong adjudication in cases that involve multiple countries. A notable report (102) suggests that Europe has faced hurdles in regulating large tech corporations due to the inefficient operations of Ireland’s Data Protection Commission, where many such companies maintain their headquarters. Specifically, an astounding 98% of GDPR cases linked to Ireland are yet to be resolved. However, EU regulatory authorities have refuted this claim and launched an official investigation (101; 103). The investigation concluded that the practices were appropriate, yet a series of improvements were suggested.
@@
Simultaneously, however, several studies indicate that DPAs are grappling with considerable resource constraints, including deficiencies in technical specialists and funding (102; 200; 51). These limitations could potentially impede the pace of enforcement advancement. Consequently, it is plausible that DPAs heavily depend on user reports to detect data rights infringements (100; 150; 109; 131; 49).
@@ 6. Towards a better data rights regime
Our review indicates that there is substantial room for enhancing data rights implementation. Although our findings focus on the GDPR, they may resonate with other similar laws, such as CCPA. Over 140 countries have comprehensive privacy laws, often drawing from the GDPR or the EU Data Protection Directive (201). Thus, our study’s implications might be applicable to various rights-based laws, raising questions about their effectiveness in truly empowering citizens. For instance, when we examine data consent, which was not within the scope of our analysis, we see that it is also currently ineffective due to reasons similar to those presented in our paper, such as users’ non-exercise of consent, inadequate implementation, and the prevalence of dark patterns (110; 120; 140; 93; 191; 193; 19; 207; 80; 55).

### v2

@@ 4. Companies’ perspective
(C6.1) Usability challenges. The process for users to exercise data rights is generally considered to have poor usability. Consider privacy policies, for instance. Numerous studies have concluded that these policies are typically too verbose and complex for the average person, as evidenced by various readability metrics, including word count and Flesch scores (136; 115; 231; 11; 166; 152; 112; 248; 193; 202; 90; 21; 235; 122). On average, these policies contain about 2000 words and are written at a 13th-grade reading level, presenting substantial comprehension challenges for many users. Consistent with this, users frequently express frustration with the difficulty of understanding privacy policies (136; 186; 19; 138). The implementation requirements according to the GDPR data rights can degrade the usability of platforms (244; 235).
@@
These usability issues, along with dark patterns that result in discouraging users’ exercise of data rights, extend beyond the right to information and impact other rights such as the rights to access, erasure, data portability, and objection (188; 129; 169; 87; 148; 90; 89; 238; 180; 236; 88; 262; 59). A platform allows users to access their data only when they email them in a specific format (236). Numerous studies also point to the difficulty of understanding downloaded data due to its volume, disorganization, use of jargon, and a lack of description (238; 183; 180; 262). Additionally, research reveals that users often struggle to find ways to delete their personal data from apps and services (188; 129; 169). For example, disproportionate amounts of time to find a “delete" (169) or “opt-out” (89) button or the need to call the service team to do so (129). Indeed, the absence of straightforward account deletion options is a pervasive issue in online services, affecting about 77% of them (87). Overall, the evidence suggests that exercising data rights is a nontrivial task for many users.
@@
(C6.2) Usability Successes. Despite the above challenges, there are instances where users have found the process of exercising their rights straightforward, particularly in the cases of the rights to access and erasure (8; 144; 238; 89). For example, a survey shows that 58% of users who had previously requested companies to erase their personal information found the process straightforward (144). Furthermore, a study found that users had no difficulty in requesting a copy of their personal information from a loyalty card provider (8). Certain websites, like Amazon, Facebook, and Google, provide online portal tools to facilitate the rights request process, eliminating the need for users to contact the companies directly (238). Consequently, the usability of exercising data rights varies significantly across different applications and data rights types.
@@
(C6.3) High-usability Data Rights Tools. Given the usability issues associated with exercising data rights, the development of user-friendly data rights tools has been a natural response. Developers and researchers proposed and launched many tools and services that allow users to easily make use of data rights (198; 138; 18; 30; 207; 189; 23; 158; 120; 119; 41; 180; 179). While several studies validate that such tools can increase the effectiveness of data rights (198; 138; 74; 18; 30; 207; 189; 41), we can also find some testimonies that raise doubts about their effects (231; 74; 200; 120; 8; 37; 180). Given this, the effects of “high usability” tools are still unclear.
@@
(C6.3.1) Effectiveness. Empirical research suggests that high-usability data rights tools can effectively assist users in asserting their data rights. Emphasis is on tools that aim to mitigate the common challenge of understanding privacy policies (198; 138; 18; 30; 207; 189; 23; 158; 120; 41; 179). They commonly employ visualization, summarization, and user engagement strategies to simplify data processing comprehension.
@@
For instance, designs using visualization techniques display icons representing collected data or diagrams illustrating data flows, aiding intuitive understanding (207; 189; 30). In contrast, some tools prioritize brevity and highlight crucial points in privacy policies, presenting users with a summarized version (138; 18; 254; 179). Users concur that these techniques–visualization and summarization–facilitate understanding of policy content compared to conventional privacy policies. Moreover, research suggests the significant role of color in enhancing comprehension, with vibrant hues being particularly effective. Such approaches could also be applied to improve the readability of data obtained via the right to access (238; 180).
@@
Another category of tools promotes user engagement as a strategy to increase awareness of policy content (120; 119). This is predicated on the established theory that active learning fosters more effective comprehension (36; 178; 243). Empirical evidence confirms enhanced user awareness when users engage actively with policy content via “Drag-and-Drop” and “Question-and-Answer” features. User experiments also reveal a higher likelihood or willingness of data rights assertion when interacting with a chatbot (41; 157). In conclusion, ample evidence supports the efficacy of designs employing visualization, summarization, or user engagement in improving the effectiveness of data rights.
@@
Interestingly, many users do not expect companies to provide easy-to-use online interfaces to exercise data rights. In an interview, while some users, particularly those more interested in their data rights, express a desire for online data rights tools, the majority view traditional means such as email or phone as the most appropriate channels to claim their access rights (8). Some users even express a preference for offline data rights exercises, favoring direct interaction with a responsible party (37; 163). Furthermore, an evaluation of a purportedly “high-usability” data dashboard (240) revealed unfavorable responses. A subset of participants demonstrated no inclination or necessity to utilize it, some voiced objections to the very notion of “high-usability” dashboards, and others dismissed data management as inconsequential.
@@
The implementation of data rights should not expose users to additional threats. Current practices, however, often fail to satisfy this criterion. Specifically, research indicates a risk of leaking personal information to unauthorized third parties when users exercise their right to access (42; 43; 63; 35). Companies must verify the identities of those making data rights requests to ensure they are the actual data owners. A flawed identification process can result in the leakage of sensitive personal information to unauthorized individuals. Several studies reveal system vulnerabilities susceptible to impersonation via counterfeit emails (42; 43; 63). These studies demonstrate how easily an attacker can deceive service providers by using an email address similar to the original one, with success rates ranging from approximately 30% to 70% (42; 43; 63). Some systems do not even implement email encryption, despite requiring users to submit identification documents via email (42; 35). Hence, severe threats often arise when users exercise their right to access.
@@
(C8.1) Efficient data use. Several studies contend that the implementation of GDPR data rights significantly enhances companies’ data processing awareness, leading to improved user data management and increased competency in utilizing data (97; 144; 65; 182; 28; 211). The growing responsibility of companies in the information economy can also foster better decision-making and risk assessment (65). Large companies, in particular, seem to strongly recognize the benefits of improved data management processes (182).
@@
However, to assess whether data rights lead to efficient data use, it is important to consider not only companies’ data processing but also users’ data sharing behaviors. If data rights cause users to be more reticent about sharing their data, it could compromise the efficiency of data use by companies, despite improvements in internal data management processes. As discussed in U4.2, the evidence on this matter is conflicting (38; 156; 45; 171; 24). While some studies suggest that empowering users may generally incentivize them to disclose more personal data (38; 156; 45; 24), one report indicates that explicitly linking user data rights with the term GDPR can have the opposite effect (171). Therefore, in contrast to previous studies, we posit that it remains uncertain whether GDPR data rights enhance data use efficiency within companies. This area warrants further research.
@@
(C8.2) Positive corporate image. There is evidence of an increased trust of users in companies that foster data rights (144; 249; 154; 95). Survey results suggest that service providers can bolster their credibility by offering options to modify/delete personal information and maintaining transparency about data processing (144; 154; 95). A separate survey indicates that U.S. consumers place greater trust in companies that voluntarily extend GDPR data rights to non-European users (249). These findings suggest that user empowerment can enhance a company’s image.
@@
However, the relationship between data rights and trust is multifaceted. The extent to which data rights can rehabilitate a company’s image, if it is initially negative, remains uncertain. Under conditions of low trust, users may question the data rights themselves (187; 19; 56; 45; 176), making it challenging for companies to anticipate the positive effects of endorsing data rights. For instance, some interviewees express skepticism about whether companies would actually delete their personal information upon request (56). Another study reveals a perception that companies deliberately obfuscate their privacy policies to discourage user engagement (19). Some users also mistrust the information presented by companies (231).
@@
Low trust in companies appears to be a widespread issue among users. Research shows that, even when users express satisfaction with privacy laws like GDPR aimed at empowering them, they tend to lack confidence in companies’ compliance with such laws (8; 144; 56; 176). Consequently, trust may not increase substantially even when companies implement data rights, given the overall low level of trust in companies.
@@
Moreover, even if data rights significantly enhance trust in companies, this could have a counter-effect: as users trust service providers more, their willingness to exercise data rights may decrease, as discussed in U2.2 (115; 198). In summary, the relationship between data rights and trust is complex and warrants further exploration.
@@ 5. Regulators’ perspective
European regulators adopt a myriad of strategies to gauge and understand the effectiveness of GDPR data rights. Primarily, they interact with citizens through the complaints they receive, serving as a valuable resource for assessing user awareness of data rights (105; 160; 116; 140; 52). The France Data Protection Authority (DPA) reported a 64% increase in complaints in 2019 compared to 2018, interpreting this as an indication that EU citizens have “strongly embraced the GDPR” (116; 140). However, as discussed in U1, this assertion may not hold true. If the volume of complaints was low prior to the enforcement of the GDPR, a 64% increase may not necessarily imply a substantial rise in user awareness and engagement with GDPR data rights.
@@
Regulators also recognize the importance of evaluating the GDPR (190; 69). Evidence suggests that regulators understand the challenges associated with implementing data rights. Among the most cited difficulties is designing GDPR-compliant ML systems (53; 103; 181). Certain regulators have also broached the challenges of implementing GDPR data rights within blockchain systems (51). They actively research ML and blockchain technology and foster discussions through conferences and seminars (53; 80; 217; 51).
Moreover, legal experts, alongside regulators, highlight the difficulties in implementing data rights, specifically the right to erasure (141; 184). In interviews, UK data protection and privacy law experts expressed mixed views on the practicality of the right to erasure (141).
They also express concerns over certain websites deciding not to collect personal information from their users post-GDPR, arguing that it degrades the online user experience and stems from a misunderstanding of the GDPR (184). However, in spite of the challenges and unintended consequences, regulators and legal professionals largely agree that the GDPR has strengthened individual empowerment (141; 69).
@@
Post-GDPR, regulators have enacted supplementary measures to enhance users’ data rights. For instance, their websites feature numerous articles and posters promoting the concept of GDPR data rights to users and companies (48; 101; 32; 49; 4).
@@
Beyond these, regulators have proposed new guidelines and devised strategies to address various issues relating to online marketing, dark patterns, and cookies (81; 102; 52). As a result, regulators have been proactive in formulating new guidelines and strategizing to enhance user data rights. However, these guidelines also underscore the substantial costs companies may incur in complying with the GDPR.
@@
Ensuring robust enforcement is paramount to the protection of data rights, encompassing the detection and penalization of corporations infringing upon these rights via fines and other sanctions. Despite substantial endeavors by regulators to mitigate the challenges encountered by users and companies (refer to R1 and R2), ample evidence underscores the current inadequacy and deficiency of GDPR data rights enforcement. (107; 212; 196; 20; 250; 58; 7; 55).
A multitude of studies examining court cases have exposed trends in GDPR enforcement (196; 20; 7; 250; 58). These analyses reveal that imposed fines have typically been marginal in relation to the comprehensive economic and societal repercussions of GDPR noncompliance. Interestingly, even DPAs have adopted a conservative stance, refraining from levying fines to their maximum potential.
@@
The uneven enforcement across Europe can significantly prolong adjudication in cases that involve multiple countries. A notable report (107) suggests that Europe has faced hurdles in regulating large tech corporations due to the inefficient operations of Ireland’s Data Protection Commission, where many such companies maintain their headquarters. Specifically, an astounding 98% of GDPR cases linked to Ireland are yet to be resolved. However, EU regulatory authorities have refuted this claim and launched an official investigation (106; 108). The investigation concluded that the practices were appropriate, yet a series of improvements were suggested.
@@
Simultaneously, however, several studies indicate that DPAs are grappling with considerable resource constraints, including deficiencies in technical specialists and funding (107; 212; 55). These limitations could potentially impede the pace of enforcement advancement. Consequently, it is plausible that DPAs heavily depend on user reports to detect data rights infringements (105; 160; 116; 140; 52).
@@ 6. Towards a better data rights regime
Our review indicates that there is substantial room for enhancing data rights implementation. Although our findings focus on the GDPR, they may resonate with other similar laws, such as CCPA. Over 140 countries have comprehensive privacy laws, often drawing from the GDPR or the EU Data Protection Directive (213). Thus, our study’s implications might be applicable to various rights-based laws, raising questions about their effectiveness in truly empowering citizens. For instance, when we examine data consent, which was not within the scope of our analysis, we see that it is also currently ineffective due to reasons similar to those presented in our paper, such as users’ non-exercise of consent, inadequate implementation, and the prevalence of dark patterns (117; 128; 149; 98; 202; 204; 22; 219; 85; 59).